### v1

Abstract
Generalizable manipulation of articulated objects remains a challenging problem in many real-world scenarios given the diverse object structures, functionalities, and goals. In these tasks, both semantic interpretations and physical plausibilities are crucial for a policy to succeed. To address this problem, we propose SAGE, a novel framework that bridges the understanding of semantic and actionable parts of articulated objects to achieve generalizable manipulation under language instructions. Given a manipulation goal specified by natural language, an instruction interpreter with Large Language Models (LLMs) first translates them into programmatic actions on the object’s semantic parts. This process also involves a scene context parser for understanding the visual inputs, which is designed to generate scene descriptions with both rich information and accurate interaction-related facts by joining the forces of generalist Visual-Language Models (VLMs) and domain-specialist part perception models. To further convert the action programs into executable policies, a part grounding module then maps the object semantic parts suggested by the instruction interpreter into so-called Generalizable Actionable Parts (GAParts). Finally, an interactive feedback module is incorporated to respond to failures, which greatly increases the robustness of the overall framework. Experiments both in simulation environments and on real robots show that our framework can handle a large variety of articulated objects with diverse language-instructed goals. We also provide a new benchmark for language-guided articulated-object manipulation in realistic scenarios. More details and demonstrations can be found at https://geometry.stanford.edu/projects/sage/.
I Introduction
Generalizable object manipulation, which involves the transfer of manipulation skills across object shapes [38, 21], object categories[14, 13, 57, 53], or even tasks [61], is gaining increasing attention in the robot learning community. A line of works has focused on learning generalizable atomic manipulation skills of articulated objects [38, 21, 14, 13, 17, 15]. Yet this remains a challenging problem, especially in real-world scenarios with complex object structures (such as home appliances) and long-horizon tasks specified by natural language instructions.
Imagine that you bought a new blender for your kitchen. You may play it with its power off, pressing its buttons, opening its lid, or taking off its container. But to make juice with it, you need to truly understand the functionalities of these parts, for example, “pressing the container turns it on, and it will mince the contents in the container”. In Gibson’s theory of distributed cognition [62], your interactions with the blender without and with prior knowledge can be explained by two different types of affordances: physical affordance and cognitive affordance. Physical affordances are for tasks that are mainly constrained by physical structures, such as the part and joint structures of articulated objects. Cognitive affordances, on the other hand, are provided by cultural conventions and involve a high-level semantic understanding of the object.
To smartly interact with varying types of articulated objects, a general framework is needed for generating and executing actions that are both semantically meaningful and physically plausible. To this end, we introduce SAGE, a framework that bridges the semantic and actionable part understandings of articulated objects to achieve generalizable manipulations under natural language instructions (Fig. LABEL:fig:teaser). Given the RGBD scan of an articulated object and a language-specified manipulation goal as input, we design an instruction interpreter with Large Language Models (LLMs) to translate the language instruction into action programs comprised of so-called action units, which are represented by 3-tuples of semantic parts, joint types, and state changes. This process also involves an image caption module applied to the RGB inputs to convert the visual contexts into language descriptions that can be understood by the LLMs. Combining large generalist Visual-Language Models (VLMs) and domain-specialist part perception models, we are able to acquire scene descriptions that are both informative and accurate about interaction-related facts. We then convert the action programs defined on the object semantic parts into executable policies through a part grounding module that maps the semantic parts into so-called Generalizable Actionable Parts (GAParts), which is a cross-category definition of parts according to their actionabilities. To achieve this, we employ a large open-vocabulary 2D segmentation model and again combine it with a domain-specialist GAPart model by merging their part proposals and ground them together to GAPart labels. From the detected GAParts, we are able to generate physically plausible part actions indicated by the pre-obtained action programs. Finally, we also introduce an interactive feedback module that actively responds to failed action steps and adjusts the overall policy accordingly, which enables the framework to act robustly under environmental ambiguities or its own failure.
We demonstrate the effectiveness of our method both in simulation environments and on real robots. Specifically, we introduce three novel benchmarks to assess crucial components of language-guided articulated-object manipulation: scene descriptions, part perceptions, and manipulation policies. Quantitatively, our method surpasses all baseline approaches. This success is attributed to our blend of a comprehensive general-purpose model and a precise domain specialist, endowing our method with both stellar performance and robust generalization capabilities. Qualitatively, we further illustrate the proficiency of our method through its performance on several demanding tasks.
To summarize, our key contributions are:
  • We introduce a novel perspective of bridging semantic and actionable parts and construct a robot system for generalizable manipulation of articulated objects.
  • We explore the idea of joining the forces of general-purpose large models and domain-specific models extensively throughout our pipeline from scene understanding, part perception, to manipulation.
  • We demonstrate our strong generalizability on a variety of different objects under diverse language instructions in both simulation environments and on real robots.
  • We also provide a new benchmark for articulated-object manipulation including evaluations of scene descriptions, part perceptions, and manipulation policies.
II Related Work
Articulated-object manipulation. Articulated object manipulation presents significant challenges due to diverse object geometries and physical characteristics. Although benchmarks by [51, 38] have been introduced, their scope remains limited. While methods like [41, 26, 4] have delved into motion planning, visual affordance learning has also gained attention [37, 55, 54, 63, 15]. Other works [7, 58] have proposed special representations for manipulation, though they are tailored mainly for suction grippers. On the other hand, [14, 13] introduce generalizable parts as a novel representation but are limited to single-part interaction and can not handle natural instructions.
Generalizable Object Manipulation. Achieving generalization in robot applications is essential yet challenging. Various works [8, 49, 18, 12, 59] employ supervised learning with motion planning to tackle generalizable tasks such as grasping. Nevertheless, these techniques often necessitate task-specific architectures and may falter in intricate manipulations. Though reinforcement learning holds promise for complex challenges [42, 2], the lack of generalizability remains unresolved [31, 16]. The ManiSkill benchmark [38] pioneers category-level object manipulation in simulations, while [44] utilizes imitation learning with RL-trained demonstrations. However, in both cases, the generalization ability is still unclear. [14, 13] first tackles generalizable manipulation in a cross-category manner, but is still limited to predefined part classes. [57, 53] have universal generalization ability but only for grasping tasks. In summary, Generalizable object manipulation is still extremely challenging.
Language models for robotics. Incorporating pre-trained language models into robot systems has been gaining tremendous attention recently. A notable segment of these works concentrates on the grounding of language to manipulation skills, essential for executing long-horizon instructions [27, 48, 36, 5, 43, 1, 17, 19, 45, 22, 60]. A hierarchical approach, seen in another subset of studies, hinges on a two-fold process: establishing a skill set and subsequently strategizing over it using large language models (LLMs) [3, 25, 47, 23, 33, 29]. However, this method is limited by the huge data collection effort, is still constrained to the predefined skill sets, and lacks visual understanding. In contrast, LLMs are employed to spawn code for low-level skills through APIs, though predicting these skills’ continuous parameters remains challenging [32]. Some employ a way-point-based method for efficiency, sidestepping low-level robot actions in favor of trajectory hand poses [46, 17]. But they also require lots of data and have no detailed planning strategy. VoxPoser [24] performs motion planning based on affordance and obstacle maps generated by LLM, but still lacks visual understanding. For articulated objects with diverse geometries, intricate structures, and physical constraints, language-guided manipulation remains a challenging and underexplored problem.
III Method
Figure 2: Context-aware instruction interpretation. With instruction and observation (RGBD image) as input, the interpreter first generates a scene description with VLM and GAPartNet[14]. Then LLM(GPT-4) takes both instruction and scene description as input and generates semantic parts and action programs. Optionally, we can input a specific user manual, and LLM will generate a target actionable part; see Sec.III-B for more details about the actionable part.
Figure 3: Scene description. To better help with action generation, the scene description should contain object information, part information, and some interaction-related information. We use expert GAPart model[10] to generate some expert descriptions as part of the VLM prompt and then generate the scene descriptions, which works well and absorb the advantages of both models
III-A Instruction interpretation
Action programs
The goal of the instruction interpreter is to translate natural-language instruction $\tilde{I}$ into programmatic instructions. We define the most basic manipulation on one single part of articulated objects as an “action unit”, which can be represented by a 3-tuple ${\bm{a}}=(\tilde{p},j,\Delta s)$ with part name $\tilde{p}$, joint type $j$, and the change of part state $\Delta s$. The part name $\tilde{p}$ is a noun in natural language which will later be sent into an open-vocabulary segmentation module; $j$ refers to the joint directly linked to part $\tilde{p}$ and indicates whether it is a revolute or a prismatic joint; $\Delta s$ is the state change of joint $j$ under the action, which is either an angle ±$\pm\theta$ if $j$ is revolute or a relative translation ±$\pm t$ w.r.t. the part bounding box if $j$ is prismatic. The final output of the interpreter should be in one of the following formats:
  • a single action unit,
  • an unordered union of multiple action units,
  • an ordered list of multiple action units,
or be a finite combination of the three formats. Following the convention of programming languages (PL), this can be viewed as a typing system with expressions:
${\bm{a}}~{}|~{}\texttt{Union}\{{\bm{a}},{\bm{a}}\}~{}|~{}\texttt{List}[{\bm{a}},{\bm{a}}]$
Here the Union expression is for non-deterministic policy generation when multiple action units can reach the same goal, e.g., both pulling the door and pressing a button result in the microwave door being opened. List is for sequential action generation when no single-step solution exists, e.g., in order to open a door, a knob must be first rotated to unlock the door.
Visual input parsing
We also incorporate a visual input parser into our instruction interpretation process, which translates the RGB inputs into language descriptions that can be understood by LLMs. A naive solution is to directly employ a large Visual-Language Model (VLM), such as Google Bard and GPT-4V, to generate scene captions. However, we notice that most of these models lack the capability to provide exact task-related facts. Inspired by [20], we recognize that the inclusion of expert facts can enhance the correctness of visual perception. Thus, we first detect the actionable parts using the GAPartNet [14] segmentation model for actionable parts and then pass it into the VLM together with the task description to generate a comprehensive scene description. Combining the expert knowledge of the specialist model and the general knowledge of VLM, we can acquire scene descriptions that are more accurate and also task-related (Fig. 3).
Language instruction interpretation
We employ GPT-4 [39], an LLM, to parse the input instruction, which outlines the intended goal. Here is an example of the overall inputs and outputs of the instruction interpreter: Certain objects, such as appliances, have very diverse and intricate structures that can hardly be covered by general knowledge. Therefore, we also accommodate the additional input of the object usage manuals, which can be easily accessed together with the purchase of the object in real-life applications. This feature allows the LLM to gain expert knowledge of the object and thereby provide more precise and relevant outputs.
III-B Part Grounding
Figure 4: Part grounding. With the observation as input, our method ensembles 2D proposals from GroundedSAM and 3D proposals from GAPartNet. then the proposals are fed to the ground as an actionable part. Leveraging ScoreNet, NMS, and PoseNet, we then present the perception results. Notice that: (1) For the part perception evaluation benchmark, we directly employ SAM[30]. However, within our manipulation pipeline, we utilize GroundedSAM, which also considers the semantic part as input. (2) As highlighted in Fig. 2, if the LLM produces a target actionable part, the grounding process is bypassed.
After getting the results from language instruction interpretation, which includes the semantic part information and corresponding semantic-level interaction strategy, we then first ground the semantic part to the physical part. In particular, with the semantic part, e.g. handle as input, we ground it to an actionable part with its pose.
Part discovery. In our methodology for part-level object perception, we leverage the strengths of both broad-spectrum 2D vision models and specialized 3D models.
Initially, we deploy the 3D GAPart Model [14] to enhance perception. This model interprets the partial RGBD point cloud $X$ to yield its set of 3D part proposals, which are masked pointclouds denoted as $\mathcal{M}_{j}^{3D}=\{M_{j}^{3D}\}$. Subsequently, our general-purpose large model processes the RGB image $I_{rgb}$ in tandem with the semantic parts $\{\tilde{p}\}$ from action unit ${\bm{a}}$, producing a series of 2D part proposals represented as $\mathcal{M}_{j}^{2D}=\{M_{j}^{2D}\}$. By synergizing the robust capabilities of general-purpose 2D vision models, such as SAM [30] and GroundedDINO [34], with the expertise of the domain-specific 3D GAPart Model, we achieve a comprehensive and nuanced part perception.
Actionable part grounding. Then we introduce an actionable part grounding model, which takes both the 2D and 3D part proposals as input and outputs the grounded GAPart labels. To this end, We first collect a dataset of actionable part features and part labels $\{{F}^{\text{dino}}_{\text{part}},l_{\text{part}}\}$. In particular, we extract a good image feature map $F^{\text{dino}}$ from DINOv2 [40]. Then, for the target actionable part denoted as $p_{j}$, we use its mask to get the part feature $F^{\text{dino}}_{j}$. We first do max-pooling to obtain the part feature and run KNN algorithm to find the grounded actionable part label denoted as $l_{j}$
$F^{\text{dino}}_{j}$ $=\text{MaxPooling}(F^{\text{dino}}[M_{j}])$
$l_{j}$ $=\text{KNN}(\{{F}^{\text{dino}}_{\text{part}},l_{\text{part}}\},F_{j}^{\text{dino}}).$
Actionable part estimation. We introduce a pose and joint estimation model to perceive grounded actionable parts. Similar to GAPartNet, the pose estimation model takes the point cloud $X$ as input and estimates the 6D part pose. Our model tackles part symmetry similar to GAPartNet, and we can directly infer the interaction joint state from the part pose thanks to the GAPart definition.
III-C Action Generation
Once we have grounded the semantic part to the actionable part, we can generate executable manipulations on this part. We first estimate the part pose as defined in GAPartNet [14]. We also compute the joint state (part axis and position) and plausible motion direction based on the joint type (prismatic or revolute). Then we generate the part actions according to these estimations. We first predict an initial gripper pose as the primary action. It is followed by a motion induced from a pre-determined strategy defined in GAPartNet [14] according to the part poses and joint states. For example, to open a door with a revolute joint, the starting position can be on the door’s rim or handle, with the trajectory being a circular arc oriented along the door’s hinge.
III-D Interactive Feedback
Algorithm 1 Interactive perception
Input initial and current part point cloud $X_{\text{part}}^{0},X_{\text{part}}^{t}$
      Output joint axis $j_{t}$ and movement $s_{t}$
1:$t_{t},R_{t}=\text{RANSAC}(\text{Umeyama}(X_{\text{part}}^{0},X_{\text{part}}^{t}))$
2:$j_{t},s_{t}=\text{GeometryInference}(R_{t},t_{t})$
Algorithm 2 Tracking and planning update
Input Gripper target and current state $\Delta g_{t}$, $\delta g_{t}$ and part target and estimated movement $\Delta s_{t}$, $s_{t}$.
1:$prompt$ = $\text{template}(\Delta g_{t},\delta g_{t},\Delta s_{t},s_{t})$triangleright Predefined Template
2:$result$ = call GPT-4 API with $prompt$
3:switch $result$ do
4:     case "continue"
5:         Continue the current strategy
6:     case "transition to next step"
7:         Finish the current action tuple and transition to the next action.
8:     case ("halt and replan")
9:         Stop the current execution and replan $\triangleright$ Some errors happened.
10:     case ("success")
11:         Success and terminate.
Up to now, we have only utilized a single initial observation $I^{0}_{rgbd}$ for generating open-loop interactions. We now introduce a mechanism to further leverage the observations acquired during the interaction process, which can update the perception results and adjust the manipulations accordingly. Toward this goal, we introduce a two-section feedback mechanism during interaction.
Interactive perception. It should be noted that occlusion and estimation errors may arise during the perception of the first observation. To address these challenges, we propose a model that leverages interactive observations to enhance operation (see Algo. 1). Consider two distinct observations, $I_{rgbd}^{0}$ and $I_{rgbd}^{t}$, captured during an interaction. A part’s motion can be detected between these two observations. We subsequently employ the mask of the moved part to deduce the actual joint state during the interaction. Applying RANSAC [9] for outlier removal and Umeyama algorithm [50], we estimate the rotation and translation between the two frames of the part’s point cloud $X_{\text{part}}^{0},X_{\text{part}}^{t}$. This computation subsequently allows us to ascertain the joint state $j_{t}$ and the current part movement states $s_{t}$.
Tracking and planning update. Throughout the interaction, tracking of both the target gripper and part states is maintained. If significant discrepancies arise, the planner has the discretion to select from one of four states (see Algo. 2): ”continue”, ”transition to the next step”, ”halt and replan”, or ”success”. For instance, if the gripper is set to rotate 60 degrees along a joint, yet the door only opens to 15 degrees, the LLM planner would opt to ”halt and re-plan”. This interaction tracking model ensures that the LLM remains cognizant of developments during interactions, permitting strategy adjustments and, where necessary, recovery from unforeseen setbacks.
Figure 5: Task examples for experiments in simulation.
Mean Opinion Score
GAPartNet
VLM
Ours
VLM
Ours
Part description $\uparrow$ 3.8 4.1 6.3 7.2 9.6
Part accuracy $\uparrow$ 8.7 4.0 6.5 6.9 9.5
Part state precision $\uparrow$ 0.0 3.7 5.6 7.7 7.8
Object&scene description $\uparrow$ 0.0 2.7 6.2 6.4 8.0
Interaction-related info.$\uparrow$ 0.3 7.3 7.6 7.6 8.7
Overall performance. $\uparrow$ 2.6 4.4 6.4 7.2 8.7
TABLE I: User study results for scene description. We solicited volunteers to assess the quality of scene descriptions. Participants were tasked with evaluating the following aspects: (1) Performance in the overall part description. (2) Accuracy in the enumeration and naming of parts. (3) Precision in part-state depiction. (4) Depiction of objects and the overall scene. (5) Description of interaction-related information. (6) Overall performance.
Figure 6: Demonstrations of real-world results. Top: ”turn on the blender” and ”turn on/off the machine”. Our method accurately understands the part semantics and actions that are not aligned. Bottom: ”open the microwave”. The mechanical structures of the microwave prevent the robot from directly pulling open the door but instead require the button to be pressed, resulting in a failure. However, our interactive feedback model can detect the failure and recognize that it should try pressing the button instead and subsequently complete the task.
Part Perception In-distribution Unseen states Unseen objects Unseen categories Accuracy AP@50 mAP AP@50 mAP AP@50 mAP AP@50 mAP PointGroup[28] 69.70 60.58 69.54 60.48 58.26 46.29 24.57 19.40 SoftGroup[52] 69.59 60.54 70.02 59.59 59.20 47.10 28.18 22.50 AutoGPart[35] 66.81 57.63 67.60 56.69 55.30 43.50 26.24 20.38 GAPartNet[14] 81.42 72.55 80.80 71.73 63.18 53.94 36.39 27.40 PartGroundedSAM [30, 40] 73.73 61.75 72.36 62.02 66.25 38.64 41.59 28.45 Ours 83.04 72.23 82.39 71.91 72.17 58.04 47.69 34.57 TABLE II: Part perception results. We have curated a novel evaluation dataset for part perception, enriched with more comprehensive data. Our method is benchmarked against 3D point cloud-based techniques, including PointGroup[28], SoftGroup[52], AutoGPart[35], and GAPartNet[14]. In alignment with the 2D branch of our approach, we also employ SAM[30] and DINOv2[40] to establish a 2D-centric baseline. For evaluation, we adopt AP@50 and mAP as our primary metrics.
Figure 7: Visual plots of the part perception results. Our method consistently achieved the highest AP@50 score for each level of generalization.
Obj. Cat. Microwave StorageFurniture Cabinet KitchenPot Remote Blender
Task ID 1 2 3 4 5 6 7 8 9 10 11 12
$*{}^{*}$[24](%) - - 13.0 - - 15.0 - - 14.2 - - -
$*{}^{*}$[14](%) 87.5 75.5 58.6 53.3 76.9 81.3 50.0 66.3 73.8 - - -
Ours(%) 98.0 96.0 94.1 83.3 80.0 95.0 79.3 89.6 83.1 85.7 60.7 42.9
TABLE III: Language-guided manipulation results. Comparison between our method with VoxPoser [24] and GAPartNet[14]. We evaluate 12 tasks with 6 different articulated objects. ”-” means the task is not implemented by the baseline. $*{}^{*}$ shares the same execution policy as our method, while $*{}^{*}$ is adapted from the unofficial version of the authors. Obj.Cat.=Object Category
Obj. Cat. # Tasks # Initial state # Target state # Instruction
Microwave 100 3 2 5
StorageFurniture 40 3 2 5
Cabinet 40 3 2 5
KitchenPot 20 1 1 5
Remote 18 1 1 5
Blender 20 1 1 5
TABLE IV: Benchmark statistics. Detailed statistics for our language-guided manipulation benchmark. For each object category, we created tasks that were randomly initialized. Obj.Cat.=Object Category
IV Experiments
IV-A Scene description.
In our instruction interpreter, a scene description is generated from the visual observations to inform the other modules of the scene context. To evaluate the quality of scene descriptions generated by our method which joins the forces of generalist and specialist models, we curate an image dataset specifically for evaluating scene descriptions for manipulation purposes. The dataset includes 145 images from 15 object categories with varying part states. We also designed six interaction-oriented metrics for assessing the qualities of the descriptions: part description, part accuracy, part state precision, object&scene description, interaction-related information, and overall performance. With these metrics, we can collect the Mean Opinion Score (MOS) through carefully designed user studies, as in [11].
We evaluate the scene descriptions generated by five methods: (1) GAPart Model: detection results are translated into natural language descriptions through a human-crafted format; (2)(3) VLM (Bard, GPT-4V), craft descriptions from the input image using Bard and GPT-4V; (4)(5) Our methods (based on Bard and GPT-4V). In the user study, participants are provided an image and the descriptions generated by the three methods at each time, and they are asked to rate the descriptions using a scale of 0-10 (with 10 being the best) according to the 6 metrics. 17 participants have been invited, and each person is asked to rate about 40 scenes on average. Table I shows the results of this user study. We found that domain-specific models can hardly provide useful object and scene-level descriptions due to their lack of general knowledge. On the other hand, VLM provides rich context but lacks accuracy in domain-related information. Our method benefits from both of them and achieves much better performances by joining their strengths.
IV-B Part perception.
We also introduce a new benchmark for part perception evaluation. In comparison with GAPartNet, our benchmark is more comprehensive and suitable for manipulation tasks. For instance, we incorporate a greater number of objects with closed parts, accounting for 12.5%, which is not included in GAPartNet[14]. To enable the evaluation of generalizations at different levels, the test data are divided into 4 subsets: in-distribution, unseen (articulation) states, unseen objects, and unseen object categories. We use the average precision AP@50 and mAP as evaluation metrics for part segmentation, which are widely adopted in prior 3D semantic/instance segmentation benchmarks such as ScanNet[6] and GAPartNet[14],
Table II presents the results of our method in comparison to various baselines for part perception. We consider GAPartNet[14], a modified version of PointGroup[28], SoftGroup[52], and AutoGPart [35] as our primary baselines. Additionally, we adapted GroundedSAM, rebranding it as PartGroundedSAM for this context. The performance metrics in Table II clearly indicate that our approach surpasses other baselines. Notably, we observed that methods based on 3D tend to under-perform on out-of-domain data, whereas 2D-centric methods yield subpar results for parts. Our methodology derives advantages from both the 2D and 3D realms, resulting in optimal performance and superior generalization capabilities.
IV-C Manipulation in Simulation Environments
We conducted our simulations using the SAPIEN environment [56] and designed 12 language-guided articulated object manipulation tasks. A comprehensive breakdown of these tasks, along with detailed statistics, is presented in Table III IV. For each category of Microwave, StorageFurniture and Cabinate, we devised 3 tasks including open from the initial open and close state and closing from the initial open state. The remaining tasks for KitchenPot, Remote and Blender are Open the lid, Press the button and Turn on the blender respectively. For each task, we carefully curated more than 5 distinct objects sourced from the GAPartNet dataset [14] that were well-suited for the respective actions. For example, we selected 5 different models of Microwaves from GAPartNet, chosen specifically for tasks like pulling the door open and pushing the door closed. Likewise, we chose 20 different StorageFurniture objects and 10 Blenders, each tailored to their corresponding tasks. To assess the robustness of our part interaction module, we conducted over 20 trials for each task. During each trial, we introduced randomization in both camera position and initial joint states, ensuring the variability of scenarios. Specifically, for the task Close the door, the initial positions of the articulated object doors were randomized within the range of $(30,60)$ degrees. To facilitate fair comparisons, we leveraged pre-trained weights from GAPartNet to identify actionable parts within the scene. Subsequently, we applied the same motion generation policy as our pipeline. The results of our experiments are summarized in Table III, showcasing the superior performance of our methods across nearly all tasks. The key drivers of our success can be attributed to our enhanced perception capabilities, which benefit from the fusion of the specialized model from GAPartNet and the generalist model of LLM.
IV-D Real-World Manipulation
In our real-world experiments, we establish an experimental setup with the UFACTORY xArm 6 and several different articulated objects for manipulation. Fig. 6 shows some featured results, and more results can be found in the supplementary video and our website. Fig. 6 top left shows a challenging case of activating a blender. The top part of the blender is perceived as a container for containing juices but functions as a button to be pressed down. Our framework effectively links its semantic and action understandings and successfully executes the task. Fig. 6 top right is an emergency stop button for robots, which requires a press (down) to halt an operation and a rotation (up) to restart it. With the auxiliary input of the user manual, our method completes both of these two tasks. Fig. 6 bottom shows a rather challenging case where the microwave door cannot be directly pulled open. Instead, it requires first pressing the button to initiate a slight door opening. In this case, our method first tries the most straightforward solution of pulling the door and fails. The interactive feedback module then detects this failure and replans a second strategy that adapts to the current environment, which finally completes the task.
V Conclusions
In this paper, we introduce a novel framework for language-guided manipulation of articulated objects. Bridging the understanding of object semantics and actionability at the part level, we are able to ground language-implied actions to executable manipulations. Throughout our framework, we also study the combination of general-purpose large vision/language models and domain-specialist models for enhancing the richness and correctness of network predictions. better handle these tasks and achieve state-of-the-art performance. We demonstrate our strong generalizability across diverse object categories and tasks. We also provide a new benchmark for language-instructed articulated-object manipulations.